# Prosoluble subgroups of the profinite completion of the fundamental group of compact 33 33-manifolds

### v2

@@ Abstract
We give a description of finitely generated prosoluble subgroups of the profinite completion of 33 33-manifold groups and virtually compact special groups.
@@ 1 Introduction
In recent years there has been a great deal of interest in detecting properties of the fundamental group $\pi_{1}(M)$ of a 33 33-manifold via its finite quotients, or more conceptually by its profinite completion. This motivates the study of the profinite completion $\widehat{\pi_{1}(M)}$ of the fundamental group of a 33 33-manifold. The study of $\widehat{\pi_{1}(M)}$ is based on the work of Henry Wilton and the second author (see [WZ17]) where it is proved that the standard splittings of 33 33-manifold groups as free product with amalgamation, HNN-extension and generally as a graph of groups (we refer to them as free constructions later) survive after profinite completion of $\pi_{1}(M)$. Namely $\widehat{\pi_{1}(M)}$ admits the (profinite) Kneser–Milnor decomposition and then the (profinite) JSJ-decomposition, and for hyperbolic manifolds with cusps (after passing to a finite sheeted cover) there is a suitable hierarchy in which the corresponding actions on profinite trees are profinitely acylindrical.
At this point one can use the profinite analogue of Bass-Serre theory for groups acting on trees. Note, however, that this theory does not have the full strength of its classical original. The main theorem of Bass-Serre theory (that describes the subgroups of free constructions) asserts that a group $G$ acting on a tree $T$ is the fundamental group of a graph of groups $(\mathcal{G},G\backslash T,D)$, where $D$ is a maximal subtree of $G\backslash T$ and $\mathcal{G}$ consists of edge and vertex stabilizers of a connected transversal of $G\backslash T$. This does not hold in the profinite case. As a consequence, subgroup theorems do not hold, in general, for profinite groups and even for free profinite products, i.e., the profinite version of the Kurosh subgroup theorem is not valid. This implies that it is reasonable to study subgroup structure of free constructions for important subclasses of profinite groups. The most important subclasses of profinite groups are prosoluble and pro-$p$ as they play the same role as finite $p$-groups and soluble groups in finite group theory. In particular, this applies to the profinite completion of 33 33-manifold groups.
Such a study for pro-$p$ subgroups was performed in [WZ18], where the structure of finitely generated pro-$p$ subgroups of profinite completions of the fundamental groups of a 33 33-manifold was described. The objective of this paper is to study prosoluble subgroups of $\widehat{\pi_{1}(M)}$.
We begin with the fundamental group of a compact hyperbolic 33 33-manifold.
@@
Let $\pi_{1}(M)$ be the fundamental group of a compact hyperbolic 33 33-manifold $M$ and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then $H$ is projective, i.e., of cohomological dimension 11 11.
@@
The proof of Theorem A uses the dramatic developments of Agol ([Ago13]), Kahn-Markovic ([KM12]) and Wise ([Wis21]), whose works implies that the fundamental groups of closed hyperbolic 33 33-manifolds are also fundamental groups of compact, virtually special cube complexes (see [AFW15] for a summary of these developments); such groups are called virtually compact special. Virtually special groups are originated in the works of Wise. These groups have a central role in modern geometric group theory and Haglund and Wise showed that many hyperbolic groups are virtually special such as hyperbolic Coxeter groups and one relator groups with torsion.
The next theorem describes torsion-free finitely generated prosoluble subgroup of virtually special groups.
@@
The next objective of this paper is to extend these results to relatively hyperbolic virtually compact special groups, in particular, to the fundamental hyperbolic 33 33-manifolds with cusps. It is based on relatively hyperbolic analogs of Wise’s Quasiconvex Hierarchy Theorem proved by Eduardo Einstein [Ein19] and the fact that it is preserved by the profinite completion established in [Zal23a].
@@
Using the [WZ18, Theorem 4.3] for 33 33-manifolds with cusps we deduce the following:
@@
Let $M$ be a hyperbolic 33 33-manifolds with cusps and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then one of the following statements holds:
@@
Finally, analyzing then $\widehat{\pi_{1}(M)}$ for Thurston’s geometries case-by-case, we obtain the classification of finitely generated prosoluble subgroups of $\widehat{\pi_{1}(M)}$ for an arbitrary compact, orientable 33 33-manifold.
@@
Let $M$ be a compact orientable 33 33-manifold. If $H$ is a finitely generated prosoluble subgroup of $\widehat{\pi_{1}(M)}$, then one of the following statements holds:
@@
    $H$ is isomorphic to a subgroup of a free prosoluble product of the pro-$p$ groups from the following list of isomorphism types:
@@
      For $p=2$: in addition to the list of (1) we have $C_{2^{m}}$; $D_{2^{k}}$; $Q_{2^{n}}$; $\mathbb{Z}_{2}\rtimes C_{2}$; the torsion-free extensions of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ by one of $C_{2}$, $C_{4}$, $C_{8}$, $D_{2}$, $D_{4}$, $D_{8}$, $Q_{16}$; the pro-22 22 extension of the Klein-bottle group $\mathbb{Z}\rtimes\mathbb{Z}$; the pro-22 22 completion of all torsion-free extensions of a soluble group $(\mathbb{Z}\rtimes\mathbb{Z})\rtimes\mathbb{Z}$ with a group of order at most 22 22;
@@
    $H$ is a subgroup of a central extension of $D_{2n}$, a tetrahedral group $T$, a octahedral group $O$ or a icosahedral group $I$ by a cyclic group of even order;
@@
    $H$ is a subgroup of the profinite completion of a 33 33-dimensional Bieberbach group $B$ (i.e, $B$ is torsion-free virtually $\mathbb{Z}^{3}$);
@@
    $H$ is a subgroup of the profinite completion of a group extension of $\mathbb{Z}^{2}\rtimes_{A}\mathbb{Z}$ by $C$ where $C$ is trivial or $C_{2}$ and $A$ is an Anosov matrix;
@@
    $H$ is an extension of a torsion-free procyclic group ^$\widehat{\mathbb{Z}}_{\sigma}$ by a subgroup $H_{0}$ of a finite free prosoluble product of finite cyclic $p$-groups with with $H_{0}$ acting either trivially on ^$\widehat{\mathbb{Z}}_{\sigma}$ or by inversion.
In this theorem we are denoting by $T$ the tetrahedral group of 1212 1212 rotational symmetries of a tetrahedron, by $O$ the octahedral group of 2424 2424 rotational symmetries of a cube or of an octahedron, by $I$ the icosahedral group of 6060 6060 rotational symmetries of a dodecahedron or an icosahedron. An Anosov matrix is a matrix $A\in\operatorname{GL}_{2}(\mathbb{Z})$ determined by an Anosov homeomorphism of the torus satisfying the following conditions $\operatorname{tr}(A)>2$ if $\det(A)=1$ and $\operatorname{tr}(A)\neq 0$ if $\det(A)=-1$ (see [GM23, Section 2]). Also, we use the notation ^$\widehat{\mathbb{Z}}_{\alpha}=\prod_{p\in\alpha}\mathbb{Z}_{p}$ where $\alpha$ is a set of primes.
@@
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group. If $H$ is a finitely generated prosoluble subgroup of $G$, then $H$ embeds into a free profinite product $\coprod_{v\in\Gamma}H_{v}$, where $H_{v}=H\cap\mathcal{G}(v)^{g}$ and $g\in G$. Moreover, if $H\neq H_{v}$ for some $v$, then either $H_{v}$ are finitely generated pro-$p$ or $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group (i.e., $C$ is cyclic, $|C|$ divides $p-1$ for any $p\in\pi$ and $[c,z]\neq 1$ for all $1\neq c\in C,0\neq z\in\widehat{\mathbb{Z}}_{\pi}$).
The celebrated Sela theorem (see [Sel97, Theorem 3.2]) states that a finitely generated group is $k$-acylindrically accessible, i.e., if it acts on a tree $T$ then $G\backslash T$ is finite and, in fact, bounded in terms of $k$ and the minimal number of generators $d(G)$. This is equivalent to the statement that $G$ has finitely many, bounded, non-trivial vertices stabilizers up to conjugation. These statements are not equivalent however for profinite groups (one can see it using Proposition 2.15). Thus the correct profinite analog of accessibility is a bound for the number of non-trivial vertex stabilizers up to conjugation. Using Theorem F we prove an analogue of Sela’s theorem for prosoluble groups.
@@
Let $G$ be a finitely generated prosoluble group acting on simply connected profinite graph $\Gamma$111simply connected means $\pi_{1}(\Gamma)=1$ and this is not the same as acyclicity in the profinite case (cf. Definition 2.4).. Then the number of vertex stabilizers of $G$, up to conjugation, is at most $d(G)$.
This paper is organized as follows. In Section 2 we introduce the basic ideas and notations of profinite Bass-Serre theory as well as prove auxiliary results. In Section 3, we study pro-$\mathcal{C}$ groups acting $k$-acylindrically on a profinite simply connected graph. Closed pro-$\mathcal{C}$ subgroups of the profinite fundamental groups of a finite graph of profinite groups with finite edge groups are studied in Section 4 and Section 5 dedicated to finitely generated prosoluble subgroups of the fundamental group of $k$-acylindrical finite graph of profinite groups. Section 6 is dedicated to prosoluble subgroups of virtually special groups. We use it in Section 7 to prove Theorem D. In Section 8 we deal with Thurston’s geometries of a compact 33 33-manifold in order to prove Theorem E.
When we deal with profinite groups our subgroups are assumed to be closed, homomorphisms to be continuous, finitely generated means topologically finitely generated. Throughout the paper we denote by $\mathcal{C}$ a variety closed to extensions according to the terminology of [RZ10]. We shall restate the results of the introduction inside of the section for easier reading.
@@ 2 Finite graphs of pro-$\mathcal{C}$ groups and its fundamental group
In this section we will introduce the basic definitions and notations of profinite Bass-Serre theory used in the paper. This section is entirely based on [Rib17].
@@ 2.1 Finite graphs of pro-$\mathcal{C}$ groups
A profinite graph is a profinite space $\Gamma$ with a distinguished non-empty subset $V(\Gamma)$ and two continuous maps $d_{i}:\Gamma\to V(\Gamma)$, $i=0,1$, which are identity on $V$.
@@
A morphism $\varphi:\Gamma\to\Delta$ of profinite graphs is a continuous map with $\varphi d_{i}=d_{i}\varphi$ for $i=0,1$.
@@
A profinite graph $\Gamma$ is connected if given any morphism of profinite graphs $\varphi$ from $\Gamma$ onto a finite graph $\Sigma$, then $\Sigma$ is connected as an abstract graph.
@@
Let $G$ be a pro-$\mathcal{C}$ group acting freely on a profinite graph $\Gamma$. The natural epimorphism $\zeta:\Gamma\to G\backslash\Gamma$ is called a Galois $\mathcal{C}$-covering of $G\backslash\Gamma$ and the associated group is denoted by $G=G(\zeta)$. If $\Gamma$ does not have non-trivial Galois $\mathcal{C}$-coverings, then $\zeta$ is called universal $\mathcal{C}$-covering and, in this case, $\pi_{1}(\Gamma)=G(\Gamma|\Delta)$ is the pro-$\mathcal{C}$ fundamental group of $\Gamma$. If $\pi_{1}(\Gamma)=1$ then we say that $\Gamma$ is $\mathcal{C}$-simply connected.
Let $\Gamma$ be a profinite graph. Define $E^{\ast}(\Gamma)=\Gamma\backslash V(\Gamma)$ and ^$\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}$ as the pro-$\mathcal{C}$ completion of $\mathbb{Z}$.
@@
$0\displaystyle 0\buildrel\over{\longrightarrow}[[\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}(E^{\ast}(\Gamma),\ast)]]\buildrel D\over{\longrightarrow}[[\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}(V(\Gamma))]]\buildrel\varepsilon\over{\longrightarrow}\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}\buildrel\over{\longrightarrow}$ 0
@@
In [Rib17, Section 3.10] we can see that `````` `every simply connected profinite graph is a profinite tree""", `````` `if $\mathcal{C}$ is the variety of all finite soluble groups, $\mathcal{C}$-simply connected and $\mathcal{C}$-tree are equivalent""".
Also, if we define
@@
to be the set of $g$-fixed points for a pro-$\mathcal{C}$ tree $T$, then $T^{g}$ is again a $\mathcal{C}$-tree (see [Rib17, Theorem 4.1.5]).
@@
Let $\Gamma$ be a pro-$\mathcal{C}$ tree. Given the vertices $v,w\in\Gamma$ we can consider $[v,w]$ as the intersection of all subtrees of $\Gamma$ containing $v$ and $w$. We call $[v,w]$ of geodesic. If $[v,w]$ is a reduced path $p_{v,w}$ from $v$ to $w$, the length $\ell([v,w])$ of $[v,w]$ is the number of edges in $p_{v,w}$.
@@
$\underline{\alpha}=(\alpha,\alpha^{\prime}):(\mathcal{G},T)\to(\mathcal{G}^{\prime},T^{\prime})$
of graphs of pro-$\mathcal{C}$ groups consists of a pair of continuous maps $\alpha:\mathcal{G}\to\mathcal{G}^{\prime}$ and $\alpha^{\prime}:T\to T^{\prime}$ such that the diagram
[…]
@@
Given a finite graph of groups with a fictitious edges $e_{1},...,e_{n}$ we can collapse this edges transforming it in a reduced finite graph of groups with the same fundamental group (see [CZ23, Remark 2.8]).

### v3

# Prosoluble subgroups of the profinite completion of the fundamental group of compact $3$-manifolds

@@ Abstract
We give a description of finitely generated prosoluble subgroups of the profinite completion of $3$-manifold groups and toral relatively hyperbolic virtually compact special groups.
@@ 1 Introduction
In recent years there has been a great deal of interest in detecting properties of the fundamental group $\pi_{1}(M)$ of a $3$-manifold and other groups of geometric nature via its finite quotients, or more conceptually by its profinite completion (see [BRW17], [BMRS20], [CW22], [Wil18]). This motivates the study of the profinite completion $\widehat{\pi_{1}(M)}$ of the fundamental group of a $3$-manifold. The study of $\widehat{\pi_{1}(M)}$ is based mainly on the work of Henry Wilton and the second author (see [WZ17]) where it is proved that the standard splittings of $3$-manifold groups as a free product with amalgamation, HNN-extension, and generally as a graph of groups (we refer to them as free constructions later) survive after profinite completion of $\pi_{1}(M)$. Namely $\widehat{\pi_{1}(M)}$ admits the (profinite) Kneser–Milnor decomposition and then the (profinite) JSJ-decomposition, and for hyperbolic manifolds with cusps (after passing to a finite sheeted cover) there is a suitable hierarchy in which the corresponding actions on profinite trees are profinitely acylindrical.
[…]
Such a study for pro-$p$ subgroups was performed in [WZ18], where the structure of finitely generated pro-$p$ subgroups of profinite completions of the fundamental groups of $3$-manifolds was described. The objective of this paper is to study prosoluble subgroups of $\widehat{\pi_{1}(M)}$.
We begin with the fundamental group of a compact hyperbolic $3$-manifold.
@@
Let $\pi_{1}(M)$ be the fundamental group of a closed hyperbolic $3$-manifold $M$ and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then $H$ is projective, i.e., of cohomological dimension $1$.
@@
The proof of Theorem A uses the dramatic developments of Agol ([Ago13]), Kahn–Markovic ([KM12]) and Wise ([Wis21]), whose works implies that the fundamental groups of closed hyperbolic $3$-manifolds are also fundamental groups of compact, virtually special cube complexes (see [AFW15] for a summary of these developments); such groups are called virtually compact special. Virtually special groups originated in the work of Wise ([HW08], [HW12]). These groups have a central role in modern geometric group theory and Haglund and Wise showed that many hyperbolic groups are virtually special such as hyperbolic Coxeter groups and one relator groups with torsion.
The next theorem describes torsion-free finitely generated prosoluble subgroups of hyperbolic virtually special groups.
@@
The next objective of this paper is to extend these results to relatively hyperbolic virtually compact special groups, in particular, to the fundamental hyperbolic $3$-manifolds with cusps. In order to prove that, our arguments use both an analogue of Wise’s Quasiconvex Hierarchy Theorem proved by Eduardo Einstein [Ein19] and the fact that it is preserved by the profinite completion established in [Zal23a].
@@
Using the [WZ18, Theorem 4.3] for $3$-manifolds with cusps we deduce the following:
@@
Let $M$ be a hyperbolic $3$-manifold with cusps and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then one of the following statements holds:
@@
Finally, analyzing $\widehat{\pi_{1}(M)}$ for Thurston’s geometries case-by-case, we obtain the classification of finitely generated prosoluble subgroups of $\widehat{\pi_{1}(M)}$ for an arbitrary compact, orientable $3$-manifold.
@@
Let $M$ be a compact orientable $3$-manifold. If $H$ is a finitely generated prosoluble subgroup of $\widehat{\pi_{1}(M)}$, then one of the following statements holds:
@@
    $H$ is isomorphic to a subgroup of a free prosoluble product of pro-$p$ groups from the following list of isomorphism types:
@@
      For $p=2$: in addition to the list of (1) we have $C_{2^{m}}$; $D_{2^{k}}$; $Q_{2^{n}}$; $\mathbb{Z}_{2}\rtimes C_{2}$; the torsion-free extensions of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ by one of $C_{2}$, $C_{4}$, $C_{8}$, $D_{2}$, $D_{4}$, $D_{8}$, $Q_{16}$; the pro-$2$ completion of the Klein-bottle group $\mathbb{Z}\rtimes\mathbb{Z}$; the pro-$2$ completion of all torsion-free extensions of a soluble group $(\mathbb{Z}\rtimes\mathbb{Z})\rtimes\mathbb{Z}$ with a group of order at most $2$;
@@
    $H$ is a subgroup of a central extension of $D_{2n}$, a tetrahedral group $T$, a octahedral group $O$ by a cyclic group of even order;
@@
    $H$ is a subgroup of the profinite completion of a $3$-dimensional Bieberbach group $B$ (i.e, $B$ is torsion-free virtually $\mathbb{Z}^{3}$);
@@
    $H$ is a subgroup of the profinite completion of a group extension of $\mathbb{Z}^{2}\rtimes_{A}\mathbb{Z}$ by $C$ where $C$ is trivial or $C_{2}$ and $A\in\operatorname{GL}_{2}(\mathbb{Z})$ (and $A$ is an Anosov matrix if $C$ is trivial), or a central extension of $\mathbb{Z}$ by $\mathbb{Z}^{2}$;
@@
    $H$ is an extension of a torsion-free procyclic group $\widehat{\mathbb{Z}}_{\sigma}$ by a subgroup $H_{0}$ of a finite free prosoluble product of finite cyclic $p$-groups with $H_{0}$ acting either trivially on $\widehat{\mathbb{Z}}_{\sigma}$ or by inversion.
In this theorem we are denoting by $T$ the tetrahedral group of $12$ rotational symmetries of a tetrahedron, by $O$ the octahedral group of $24$ rotational symmetries of a cube or of an octahedron. An Anosov matrix is a matrix $A\in\operatorname{GL}_{2}(\mathbb{Z})$ determined by an Anosov homeomorphism of the torus satisfying the following conditions $\operatorname{tr}(A)>2$ if $\det(A)=1$ and $\operatorname{tr}(A)\neq 0$ if $\det(A)=-1$ (see [GM23, Section 2]). We use the notation $\widehat{\mathbb{Z}}_{\alpha}=\prod_{p\in\alpha}\mathbb{Z}_{p}$ where $\alpha$ is a set of primes. Also, recall that a profinite Frobenius group $\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a group such that $C$ is cyclic, $|C|$ divides $p-1$ for any $p\in\pi$ and $[c,z]\neq 1$ for all $1\neq c\in C,0\neq z\in\widehat{\mathbb{Z}}_{\pi}$.
@@
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group. If $H$ is a finitely generated prosoluble subgroup of $G$, then $H$ embeds into a free profinite product $\coprod_{v\in\Gamma}H_{v}$, where $H_{v}=H\cap\mathcal{G}(v)^{g}$ and $g\in G$. Moreover, if $H\neq H_{v}$ for some $v$, then either $H_{v}$ is finitely generated pro-$p$ or $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group.
The celebrated Sela theorem (see [Sel97, Theorem 3.2]) states that a finitely generated group $G$ is $k$-acylindrically accessible, i.e., if $G$ is the fundamental group of a reduced $k$-acylindrical graph of groups $(\mathcal{G},\Gamma)$, then $\Gamma$ is finite and, in fact, bounded in terms of $k$ and the minimal number of generators $d(G)$. This is equivalent to the statement that $G$ acts on the corresponding Bass-Serre tree with bounded number of maximal vertex stabilizers up to conjugation. These statements are not equivalent, however, for profinite groups (one can see it using Proposition 2.15). We suggest that the correct profinite analog of accessibility is a bound for the number of maximal vertex stabilizers up to conjugation. Using Theorem F we prove an analogue of Sela’s theorem for prosoluble groups.
@@
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups and $G$ its profinite fundamental group. If $H$ is a finitely generated prosoluble subgroup of $G$, then the number of maximal vertex stabilizers of $H$ acting on the standard profinite Bass-Serre tree, up to conjugation, is at most $d(H)$.
This paper is organized as follows. In Section 2 we introduce the basic ideas and notations of the profinite Bass-Serre theory as well as prove auxiliary results. In Section 3, we study pro-$\mathcal{C}$ groups acting $k$-acylindrically on a profinite simply connected graph. Closed pro-$\mathcal{C}$ subgroups of the profinite fundamental groups of a finite graph of profinite groups are studied in Section 4, where we prove Theorem G. We prove Theorem F in Section 5. Section 6 is dedicated to prosoluble subgroups of virtually special groups, where Theorems A, B, C and D are proved. We aplly them in Section 7 to the profinite completion of arithmetic hyperbolic manifold groups. In Section 8 we deal with Thurston’s geometries of a compact $3$-manifold in order to prove Theorem E.
When we deal with profinite groups our subgroups are assumed to be closed, homomorphisms to be continuous, finitely generated means topologically finitely generated. Throughout the paper we denote by $\mathcal{C}$ a class of finite groups closed for subgroups, quotients and extensions. We shall restate the results of the introduction inside of the section for easier reading.
@@ 2 Finite graphs of pro-$\mathcal{C}$ groups and its fundamental group
In this section we will introduce the basic definitions and notations for profinite Bass-Serre theory used in this paper. This section is entirely based on [Rib17].
@@ 2.1 Finite graphs of pro-$\mathcal{C}$ groups
A profinite graph is a profinite space $\Gamma$ with a distinguished non-empty subset $V(\Gamma)$ and two continuous maps $d_{i}:\Gamma\to V(\Gamma)$, $i=0,1$, which are the identity on $V(\Gamma)$.
@@
A morphism $\varphi:\Gamma\to\Delta$ of profinite graphs is a continuous map such that $\varphi d_{i}=d_{i}\varphi$ for $i=0,1$.
@@
A profinite graph $\Gamma$ is connected if every finite quotient graph of $\Gamma$ is connected as an abstract graph.
@@
Let $G$ be a pro-$\mathcal{C}$ group acting freely on a profinite graph $\Gamma$. The natural epimorphism $\zeta:\Gamma\to G\backslash\Gamma$ is called a Galois $\mathcal{C}$-covering of $\Delta=G\backslash\Gamma$ and the associated group is denoted by $G=G(\zeta)$. If $\Gamma$ does not have non-trivial Galois $\mathcal{C}$-coverings, then $\zeta$ is called universal $\mathcal{C}$-covering and, in this case, $\pi_{1}(\Gamma)=G(\Gamma|\Delta)$ is the pro-$\mathcal{C}$ fundamental group of $\Gamma$. If $\pi_{1}(\Gamma)=1$ then we say that $\Gamma$ is $\mathcal{C}$-simply connected.
Let $\Gamma$ be a profinite graph. Define $E^{\ast}(\Gamma)=\Gamma\backslash V(\Gamma)$ and $\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}$ as the pro-$\mathcal{C}$ completion of $\mathbb{Z}$.
@@
$0\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits}[[\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}(E^{\ast}(\Gamma),\ast)]]\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{D}}[[\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}(V(\Gamma))]]\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{\varepsilon}}\widehat{\mathbb{Z}}_{\widehat{\mathcal{C}}}\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits}0$
@@
[…]
We say that a profinite group $G$ acts on the profinite tree $T$ if it acts continuously on $T$ and the action commutes with $d_{0}$ and $d_{1}$. We denote by $G_{t}$ the stabilizer of $t\in T$ in $G$.
We define
@@
to be the set of $g$-fixed points for a pro-$\mathcal{C}$ tree $T$, then $T^{g}$ is again a $\mathcal{C}$-tree (see [Rib17, Theorem 4.1.5]). The subgroup of $G$ generated by all vertex stabilizers usually will be denoted by $\widetilde{G}$.
@@
Let $\Gamma$ be a pro-$\mathcal{C}$ tree. Given the vertices $v,w\in\Gamma$ we can consider $[v,w]$ as the intersection of all pro-$\mathcal{C}$ subtrees of $\Gamma$ containing $v$ and $w$. We call $[v,w]$ a geodesic. The length $\ell([v,w])$ of $[v,w]$ is the number of edges in $[v,w]$.
@@
$\underline{\alpha}=(\alpha,\alpha^{\prime}):(\mathcal{G},\Gamma)\to(\mathcal{G}^{\prime},\Gamma^{\prime})$
of graphs of pro-$\mathcal{C}$ groups consists of a pair of continuous maps $\alpha:\mathcal{G}\to\mathcal{G}^{\prime}$ and $\alpha^{\prime}:\Gamma\to\Gamma^{\prime}$ such that the diagram
$\mathcal{G}$$\mathcal{G}^{\prime}$$\Gamma$$\Gamma^{\prime}$$\alpha$$\alpha^{\prime}$$\pi$$\pi^{\prime}$
@@
Given a finite graph of groups with fictitious edges $e_{1},...,e_{n}$ we can collapse this edges transforming it in a reduced finite graph of groups with the same fundamental group (see [CZ23, Remark 2.8]).